# Computable Thermodynamic Formalism

### v1

@@ 5.1. Jacobian and the transfer operator
Since $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$, the inequalities in (5.19) must be equalities, which implies $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. This implies $\mu\in\mathcal{E}_{0}(T,\phi)$ (recall (5.17)).
We now establish the backward implication. Suppose $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. By (3.2) and (5.18), we obtain $\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle=h_{\mu}(T)=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}=\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we have $1\geqslant\mu\bigl(T^{i}(Y)\bigr)\geqslant\mu(Y)=1$ for each $i\in\mathbb{N}_{0}$. Thus $\mu\bigl(\bigcap_{i\in\mathbb{N}_{0}}T^{i}(Y)\bigr)=1$. Since $J$ satisfies property (ii) in Theorem 5.6, by Theorem 5.5, $J$ is a Jacobian on $Y$ for $T$ with respect to $\mu$. By (1.2), we conclude that $\mu\in\mathcal{M}(X,T;Y,J)$. ∎
@@
Assume that $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions $J_{n}\colon X\rightarrow[0,+\infty)$ with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$ such that $J_{n}$ is nonnegative on $Y_{n}$ and Borel for each $n\in\mathbb{N}$. Then $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
@@
By Theorem 5.7, $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $\bigl(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}}\bigr)$. Since $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact, by Proposition 3.21 (i), $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact. Note that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}=\{\mu_{n}\}$ for each $n\in\mathbb{N}$ by (1.3). Therefore, by Proposition 3.20 (i), $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
We devote the rest of this subsection to a proof of Theorem 5.7. The proof proceeds by constructing a sequence of uniformly lower semi-computable open sets and expressing the set $\mathcal{M}(X,T_{n};Y_{n},J_{n})$ as the complement of their union. The uniform recursive compactness then follows from established properties of recursively compact spaces.
@@
    For each $x\in X$, $\{h_{m,n,k}(x)\}_{m\in\mathbb{N}}$ is non-decreasing and $h_{m,n,k}(x)\to\mathbbm{1}_{Y_{n,k}}(x)$ as $m\to+\infty$.
@@
By the uniform computability of $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$, Proposition 3.23 implies that $\bigl\{\tau_{s}^{+}\cdot h_{m,n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly computable functions. Since $\{Q_{v}\}_{v\in\mathbb{N}}$ is uniformly lower semi-computable open in $\bigl(\mathbb{R},\,d_{\mathbb{R}},\,\mathcal{S}_{\mathbb{Q}}\bigr)$, by Proposition 3.14, $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$. Note that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Then since $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, by Proposition 3.20 (iii), $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Note that by property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ and the hypotheses of Theorem 5.7, we have that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\subseteq Y_{n,k}\subseteq X_{n}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Hence, since $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions with respect to $\{X_{n}\}_{n\in\mathbb{N}}$, by Proposition 3.20 (v), we obtain that $\bigl\{T_{n}\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr)\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Since $X$ is recursively compact, and $\mathcal{S}_{\mathbb{Q}}=\{q_{v}\}_{v\in\mathbb{N}}$, by (5.23), $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Thus, since $\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$, by Proposition 3.20 (ii), the sequence $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open. Hence, it follows from Proposition 3.16 that $\bigl\{W_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly upper semi-computable functions.
@@
We next show that $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions. Denote $R_{v}\coloneqq(q_{v},+\infty)$ for each $v\in\mathbb{N}$. Since $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ are two sequences of uniformly computable nonnegative functions, and $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$, by (5.21), Definition 3.15, and Proposition 3.23, $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$. Hence, by Proposition 3.16, there exists a sequence $\bigl\{Y_{m,s,v}^{n,k}\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ of uniformly lower semi-computable open sets such that $\bigl(V_{m,s}^{n,k}\bigr)^{-1}(R_{v})\cap Y_{n}=Y_{m,s,v}^{n,k}\cap Y_{n}$ for all $m,\,s,\,v\in\mathbb{N}$ and $(n,k)\in L$. Indeed, it is not hard to derive from (5.21) and property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ that
@@
$\int\!W_{m,s}^{n,k}\,\mathrm{d}\mu$ $=\int_{T_{n}(Y_{n,k})}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)^{-1}\,\mathrm{d}\mu=\int_{Y_{n,k}}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\,\mathrm{d}\bigl(\mu\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)\bigr)$
@@
In this subsection, we demonstrate how to apply Theorem 1.3 to different systems by constructing appropriate sequences $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$, with the main applications being Theorems 5.9 and 5.10.
@@
    Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system. Assume that $\{Y_{n}\}_{n\in\mathbb{N}}$ is a sequence of open sets and $\{J_{n}\}$ is a sequence of Borel functions satisfying the hypotheses of Theorem 5.7.
@@
    $\{C_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(X,\,\rho,\,\mathcal{S})$ such that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ and $\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$.
@@
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a Borel-measurable transformation, and $\mu\in\mathcal{M}(X,T)$. Assume that $x$ is a point in $X$ with $\mu(\{x\})>0$. Then $x$ is a periodic point of $T$. If we assume in addition that $\mu$ is ergodic, then $\mu=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{T^{i}(x)}$, where $n$ is the period of $x$.
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions. Employing Theorem 5.9, we can establish the computability of equilibrium states for some expanding Thurston maps and Hölder continuous potentials (see e.g.  Theorem 1.4).
@@
Next, we establish the following result.
@@
Since $\{Y_{n}\}_{n\in\mathbb{N}}$ satisfies the hypotheses of Theorem 1.3, there exist two recursively enumerable sets K,LK,\,L with $L\subseteq\mathbb{N}\times K$, and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of open sets such that $Y_{n,k}$ is admissible for $T_{n}$ for each $(n,k)\in L$. Since $L$ is a recursively enumerable set, there exists a function $f\colon\mathbb{N}\rightarrow L_{n}$ with $f(\mathbb{N})=L_{n}$. Define $Y_{m}^{\prime}\coloneqq Y_{f(m)}\smallsetminus\bigcup_{k=1}^{m-1}Y_{f(k)}$ for each $m\in\mathbb{N}$. Then by Definition 5.2, $(X,\,\rho,\,T_{n},\,Y_{n},\,\{Y^{\prime}_{m}\}_{m\in\mathbb{N}},\,\mu)$ is admissible. Note that $\phi_{n}$ is a bounded Borel function by item (ii) of the Assumptions. Then we have $\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}\in\mathbb{R}$ for each $\mu\in\mathcal{P}(X)$ and each $n\in\mathbb{N}$. Hence, by item (iii) of the Assumptions and Theorem 5.6, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $Y_{n}\subseteq X_{n}$ by item (ii) of the Assumptions. Then by (1.1), we have $\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Thus, by item (iv) of the Assumptions, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$.

### v2

@@ 5.1. Jacobian and the transfer operator
Since $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$, the inequalities in (5.18) must be equalities, which implies $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. This implies $\mu\in\mathcal{E}_{0}(T,\phi;Y)$ (recall (5.16)).
We now establish the backward implication. Suppose $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. By (3.2) and (5.17), we obtain $\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle=h_{\mu}(T)=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}=\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we have $1\geqslant\mu\bigl(T^{i}(Y)\bigr)\geqslant\mu(Y)=1$ for each $i\in\mathbb{N}_{0}$. Thus $\mu\bigl(\bigcap_{i\in\mathbb{N}_{0}}T^{i}(Y)\bigr)=1$. Since $J$ satisfies property (ii), by Theorem 5.5, $J$ is a Jacobian on $Y$ for $T$ with respect to $\mu$. By (1.2), we conclude that $\mu\in\mathcal{M}(X,T;Y,J)$. ∎
Finally, we state the following lemma to weaken property (iii) in Definition 5.2.
Lemma 5.7.
Let $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\mu)$ satisfy properties (i), (ii), (iv), and (v) of Definition 5.2. Assume that $Y_{k}$ is admissible for $T$ for each $k\in\mathbb{N}$. Then there exists a sequence $\{Y_{k}^{\prime}\}_{k\in\mathbb{N}}$ of Borel subsets such that $(X,\,\rho,\,T,\,Y,\,\{Y_{k}^{\prime}\}_{k\in\mathbb{N}},\,\mu)$ is an admissible sextuple.
Proof.
Set $Y_{k}^{\prime}\coloneqq Y_{k}\smallsetminus\bigcup_{n=1}^{k-1}Y_{n}$ for each $k\in\mathbb{N}$. Then we have $\bigcup_{k\in\mathbb{N}}Y_{k}^{\prime}=\bigcup_{k\in\mathbb{N}}Y_{k}=Y$. By Definition 5.1, it follows from $Y_{k}^{\prime}\subseteq Y_{k}$ that $Y_{k}^{\prime}$ is admissible for $T$ for each $k\in\mathbb{N}$. Hence, by definition, $\{Y_{k}^{\prime}\}_{k\in\mathbb{N}}$ is a sequence of pairwise disjoint admissible sets for $T$, which implies that property (iii) of Definition 5.2 is satisfied for the sextuple $(X,\,\rho,\,T,\,Y,\,\{Y_{k}^{\prime}\}_{k\in\mathbb{N}},\,\mu)$. Therefore, by Definition 5.2, $(X,\,\rho,\,T,\,Y,\,\{Y_{k}^{\prime}\}_{k\in\mathbb{N}},\,\mu)$ is admissible. ∎
@@
Suppose $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions $J_{n}\colon X\rightarrow[0,+\infty)$ with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$ such that $J_{n}$ is nonnegative on $Y_{n}$ and Borel for each $n\in\mathbb{N}$. Then $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
@@
By Theorem 5.8, $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. By Proposition 3.21 (i), since $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact, so is $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\}_{n\in\mathbb{N}}$. Note that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}=\{\mu_{n}\}$ for each $n\in\mathbb{N}$ by (1.3). Therefore, by Proposition 3.20 (i), $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
We devote the rest of this subsection to a proof of Theorem 5.8. Our proof proceeds by constructing a sequence of uniformly lower semi-computable open sets and expressing the set $\mathcal{M}(X,T_{n};Y_{n},J_{n})$ as the complement of their union. The uniform recursive compactness then follows from established properties of recursively compact spaces.
@@
    For each $x\in X$, $\{h_{m,n,k}(x)\}_{m\in\mathbb{N}}$ is nondecreasing and $h_{m,n,k}(x)\to\mathbbm{1}_{Y_{n,k}}(x)$ as $m\to+\infty$.
@@
By the uniform computability of $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$, Proposition 3.23 implies that $\bigl\{\tau_{s}^{+}\cdot h_{m,n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly computable functions. Since $\{Q_{v}\}_{v\in\mathbb{N}}$ is uniformly lower semi-computable open in $\bigl(\mathbb{R},\,d_{\mathbb{R}},\,\mathcal{S}_{\mathbb{Q}}\bigr)$, by Proposition 3.14, the sequence $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$. Note that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Then since $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, by Proposition 3.20 (iii), $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Note that by property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ and the hypotheses of Theorem 5.8, we have that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\subseteq Y_{n,k}\subseteq X_{n}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Hence, since $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions with respect to $\{X_{n}\}_{n\in\mathbb{N}}$, by Proposition 3.20 (v), we obtain that $\bigl\{T_{n}\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr)\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Since $X$ is recursively compact, and $\mathcal{S}_{\mathbb{Q}}=\{q_{v}\}_{v\in\mathbb{N}}$, by (5.22), $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Thus, since $\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$, by Proposition 3.20 (ii), the sequence $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open. Hence, it follows from the implication from (ii) to (i) in Proposition 3.16 that $\bigl\{W_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly upper semi-computable functions.
@@
We next show that $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions. Denote $R_{v}\coloneqq(q_{v},+\infty)$ for each $v\in\mathbb{N}$. Since $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ are sequences of uniformly computable nonnegative functions, and $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$, by (5.20), Definition 3.15, and Proposition 3.23, $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$. Hence, by Proposition 3.16, there exists a sequence $\bigl\{Y_{m,s,v}^{n,k}\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ of uniformly lower semi-computable open sets such that $\bigl(V_{m,s}^{n,k}\bigr)^{-1}(R_{v})\cap Y_{n}=Y_{m,s,v}^{n,k}\cap Y_{n}$ for all $m,\,s,\,v\in\mathbb{N}$ and $(n,k)\in L$. Indeed, it is not hard to derive from (5.20) and property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ that
@@
$\int\!W_{m,s}^{n,k}\,\mathrm{d}\mu$ $\geqslant\int_{T_{n}(Y_{n,k})}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)^{-1}\,\mathrm{d}\mu=\int_{Y_{n,k}}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\,\mathrm{d}\bigl(\mu\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)\bigr)$
@@
In this subsection, we demonstrate how to apply Theorem 1.3 to different systems by constructing appropriate sequences $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$, with the main applications being Theorems 5.10 and 5.11.
@@
    Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system. Let $\{Y_{n}\}_{n\in\mathbb{N}}$ be a sequence of open sets and $\{J_{n}\}_{n\in\mathbb{N}}$ be a sequence of Borel functions that satisfy the hypotheses of Theorem 5.8.
@@
    $\{C_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly recursively compact subsets in $(X,\,\rho,\,\mathcal{S})$ such that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ and $\mathcal{E}_{0}(T_{n},\phi_{n};Y_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$. [^6] For the definition of the set $\mathcal{E}_{0}(T_{n},\phi_{n};Y_{n})$, see (5.16); its existence and uniqueness are discussed immediately thereafter.
@@
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a Borel-measurable transformation, and $\mu\in\mathcal{M}(X,T)$. Suppose $x\in X$ satisfies $\mu(\{x\})>0$. Then $x$ is a periodic point of $T$. If we assume in addition that $\mu$ is ergodic, then $\mu=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{T^{i}(x)}$, where $n$ is the period of $x$.
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions. Before detailing the technical statements, we emphasize their applicability to expanding Thurston maps and rational maps. Specifically, Theorem 5.10 can be applied to establish the computability of equilibrium states for expanding Thurston maps with Hölder continuous potentials. As a corollary, Theorem 5.11 directly implies the computability of equilibrium states for rational maps with hyperbolic Hölder potentials.
However, a full treatment of general potentials requires verifying the computability of the associated Jacobians, a task that relies on independent techniques such as the cone method introduced in [BHLZ25, Subsection 4.1]. To demonstrate the efficacy of our approach while avoiding the extensive technicalities associated with these estimates, we restrict our explicit application in this article to expanding Thurston maps with the zero potential (see Theorem 1.4).
@@
Next, we establish the following claim.
@@
Since $\{Y_{n}\}_{n\in\mathbb{N}}$ satisfies the hypotheses of Theorem 1.3, there exist two recursively enumerable sets K,LK,\,L with $L\subseteq\mathbb{N}\times K$, and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of open sets such that $Y_{n,k}$ is admissible for $T_{n}$ for each $(n,k)\in L$. Since $L$ is a recursively enumerable set, there exists a function $f\colon\mathbb{N}\rightarrow L_{n}$ with $f(\mathbb{N})=L_{n}$. Define $Y_{m}^{\prime}\coloneqq Y_{f(m)}\smallsetminus\bigcup_{k=1}^{m-1}Y_{f(k)}$ for each $m\in\mathbb{N}$. Then by Definition 5.2, $(X,\,\rho,\,T_{n},\,Y_{n},\,\{Y^{\prime}_{m}\}_{m\in\mathbb{N}},\,\mu)$ is admissible. Note that $\phi_{n}$ is a bounded Borel function by item (ii) in the Assumptions. Then we have $\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}\in\mathbb{R}$ for each $\mu\in\mathcal{P}(X)$ and each $n\in\mathbb{N}$. Hence, by item (iii) in the Assumptions and Theorem 5.6, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n};Y_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $Y_{n}\subseteq X_{n}$ by item (ii) in the Assumptions. Then by (1.1), we have $\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Thus, by item (iv) in the Assumptions, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n};Y_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$.